An AP has 21 terms.The sum of 10th,11th, and 12th terms is 129 and the sum of the last 3 terms is 237. Find the AP
Answer by richwmiller(17219)

a+10d+a+11d+a+12d=129,
a+19d+a+20d+a+21d=237
a = -1, d = 4
t=-1+4n